Make up a sample of size n = 6 for which the average is −7 (note the (−) sign). Not all 6 numbers should be identical.
Answer #1

All six members are identical.

Given, n = 6 , Mean = -7

Mean = X / n

-7 = X / 7

X = -49

Therefore,

Value of single sample = -49 / 7 = -7

Sample = -7 , -7 , -7 , -7 , -7 , -7

[ Here since all members are identical, value of single member = average = -7 ]